OpenStudy (anonymous):

what is sin 2pi/3? how do you solve wihtout a calculator?

OpenStudy (anonymous):

Root3 / 2.

OpenStudy (anonymous):

\[\sqrt{3}\div2\]

OpenStudy (anonymous):

unit circle

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/cheat_table.aspx look at unit circle on last page of trig cheat sheet. locate \[\frac{2\pi}{3}\] on the unit circle. the second coordinate is sine

myininaya (myininaya):

i will show you how to do this one without a calculator :)

myininaya (myininaya):

pretend we have an equilateral

myininaya (myininaya):